OpenStudy (anonymous):

Please help me!....The equation -16^(2)+25t models the height of a soccer ball t seconds after it is kicked. when will the height of the ball reach 9 feet?

OpenStudy (kropot72):

Is this the equation? \[-16^{2}+25t\]

OpenStudy (anonymous):

please look at my question i'll give you a medal if u answer please

OpenStudy (anonymous):

yeah, can you please help me?

OpenStudy (anonymous):

wish i could i remember learning that but don't remember how to do it sorry! maybe replace t with 9

OpenStudy (kropot72):

\[h=-(16)^{2}+25t\] \[9=-256+25t\] 25t = 256 + 9 = 265 t = 265/25 = 9 seconds